# Jaime Fernández Fisac and Chen Lv

### v2

@@ Abstract
where $\mu_{\theta}$ calculates the posterior mean of the noise at $k-1$ step from $\tilde{\mathbf{u}}_{k}$ and DDPM denoiser output $\mathcal{D}_{\theta}(\cdot)$, and $\sigma_{k}$ is the standard deviation according to the fixed noise schedule.
@@
3.4 3.4 Controllable and Compositional Generation
In many applications, we want to generate scenarios that satisfy a specific user requirement $y$ without retraining the model. For example, $y$ can be defined as the goal or the reference path for individual agents, or it can describe a soft constraint, such as obeying the speed limit or avoiding collisions. From the perspective of optimal control (Equation 1), we modify the optimization objective to: $\mathcal{J}_{\theta}({\mathbf{x}},\mathbf{u};\mathbf{c})+\mathcal{J}_{y}({\mathbf{x}},\mathbf{u};\mathbf{c})$. Plugging into the EBM representation, we obtain a new conditional distribution: $p(\mathbf{u}|\mathbf{c},y)\propto p(\mathbf{u}|\mathbf{c})p(y|\mathbf{u},\mathbf{c})$, where $p(\mathbf{u}|\mathbf{c})$ is the data distribution we approximated through generative modeling and $p(y|\mathbf{u},\mathbf{c})$ is the likelihood of $y$. This immediately resembles the compositionality in EBM (Du and Mordatch, 2019) and Classifier Guidance in diffusion model (Dhariwal and Nichol, 2021). Specifically, we can sample the reverse process with a conditional score function:
@@
where $p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})$ is the likelihood of $y$ given the noised action $\tilde{\mathbf{u}}$ at step $k$. It is important to note that $p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})$ is not equivalent to the likelihood of $y$ in the data distribution $p(y|\mathbf{u},\mathbf{c})$, therefore it is typically required to train a separate model (Dhariwal and Nichol, 2021; Janner et al., 2022). However, we can utilize practical approximation to the gradient of noised likelihood with the gradient of $\mathcal{J}_{y}$ for guidance, which enables flexible composition and controllability with additional objectives without training.
In the reverse diffusion process, the guidance is implemented by subtly altering the predicted mean of the model at each denoising step (Zhong et al., 2023b, a). Initially, we can directly approximate $\log p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})\approx\log p(y|\tilde{\mathbf{u}},\mathbf{c})=\mathcal{J}_{y}({\mathbf{x}}(\tilde{\mathbf{u}}),\tilde{\mathbf{u}};\mathbf{c})$; therefore, the denoising mean is modified to impose guidance in the form of a score function $\mathcal{J}_{y}$ as:
@@
where $\lambda$ is a parameter that controls the strength of the guidance.
However, calculating the gradient based on the mean of noisy actions is difficult and inaccurate, and thus manipulating the noise mean using the noisy gradient can result in errors and instability. To address this, we utilize an alternative approach (Jiang et al., 2023) to approximate $\log p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})\approx\mathcal{J}_{y}({\mathbf{x}}(\mathcal{D}_{\theta}(\tilde{\mathbf{u}})),\mathcal{D}_{\theta}(\tilde{\mathbf{u}});\mathbf{c})$. This means calculating the objective function using the one-step generation result from the denoiser rather than the noisy mean. The modification in the denoising step is expressed as:
@@
where the gradient $\nabla_{\tilde{\mathbf{u}}_{k}}$ is calculated with respect to the noisy actions $\tilde{\mathbf{u}}_{k}$ and necessitates differentiation through the denoiser $\mathcal{D}_{\theta}$.
4 4. Versatile Behavior Diffusion
Figure 3: Neural network structure of the proposed VBD model. The input scenario tokens are encoded through a query-centric Transformer scene context encoder. The behavior predictor generates marginal multi-modal trajectories. The denoiser predicts the joint multi-agent future trajectories while attending to themselves and the condition tokens. During inference, the predicted behavior priors or user-defined model-based objectives $\mathcal{J}$ can be used to guide the denoising process to generate desired scenarios.
4.1 4.1 Model Structure
@@
The Versatile Behavior Diffusion (VBD) model consists of three main components, as illustrated in Figure 3. The scene encoder $\mathcal{E}_{\phi}:\mathbf{c}\mapsto\hat{\mathbf{c}}$ encodes the scene context $\mathbf{c}$ into its latent representation $\hat{\mathbf{c}}$ using query-centric attention Transformers (Shi et al., 2023). Leveraging rich scene context information from encoder, the denoiser $\mathcal{D}_{\theta}:(\hat{\mathbf{c}},\tilde{\mathbf{u}},k)\mapsto\hat{\mathbf{u}}$ directly predict a joint clean control sequence $\hat{\mathbf{u}}$ from $\hat{\mathbf{c}}$ and noised control $\tilde{\mathbf{u}}$ at step $k$. The behavior predictor $\mathcal{P}_{\psi}:(\hat{\mathbf{c}},\{\zeta^{i}\}_{i=1}^{M})\mapsto\{\text{Cat}_{M}(\hat{\mathbf{u}}^{a},\hat{\omega}^{a})\}_{a=1}^{A}$ predicts an $M$-mode marginal categorical trajectory distribution of each agent from $\hat{\mathbf{c}}$ with the help of a set of representative static end-point anchors $\{\zeta^{i}\}_{i=1}^{M}$ extracted from data (Shi et al., 2023). All three modules utilize a stack of query-centric self-attention and cross-attention blocks for flexibility and scalability. Details regarding the model architecture are illustrated below.
@@
Scene Context. The scene conditions are divided into three categories: agents $\mathbf{c}_{a}\in\mathbb{R}^{A\times T_{h}\times D_{a}}$, map polylines $\mathbf{c}_{m,pl}\in\mathbb{R}^{M_{l}\times M_{p}\times D_{p}}$ and traffic lights $\mathbf{c}_{m,tl}\in\mathbb{R}^{M_{t}\times D_{t}}$. Here, $T_{h}$ denotes the number of historical steps, $M_{l}$ the number of polylines, $M_{p}$ the number of waypoints per polyline, $M_{t}$ the number of traffic lights, and $M_{l}+M_{t}=M$ represents the combined count of map elements. The feature sizes for agents, polylines, and traffic lights are represented by $D_{a}$, $D_{p}$, and $D_{t}$, respectively. The agent history tensor records each agent’s historical state, including $x,y$ coordinates, heading angle ($\psi$), velocities ($v_{x},v_{y}$), and bounding box dimensions ($l,w,h$), along with the agent type. Each map polyline, comprising $M_{p}$ waypoints, includes attributes like $x,y$ coordinates, direction angle, the traffic light state controlling the lane, and lane type. The traffic lights tensor encompasses $x,y$ coordinates of stop points and the state of each traffic light. Before encoding, positional attributes of all elements are converted into their local coordinate systems; for agents, the reference point is their last recorded state, and for map polylines, it is the location of the first waypoint.
@@
Scene Encoder. We encode the agent history tensor utilizing a shared GRU network, which is then combined with the agent type embedding. For map polylines, an MLP is employed for encoding, followed by max-pooling along the waypoint axis; for traffic lights, we only encode their light status using an MLP. These tensors are then concatenated to form the initial scene encoding. The initial scene encoding is further processed using $L_{\mathcal{E}}$ query-centric Transformer layers to symmetrically encode the interrelationships among scene components, resulting in a uniform dimension for scene features $\hat{\mathbf{c}}\in\mathbb{R}^{(A+M)\times D}$. In this approach, each scene element is translated into its local coordinate system and encoded with query-centric features, and the relative position of each pair of scene elements is calculated and encoded as edge attributes.
@@
Denoiser. The denoiser $\mathcal{D}_{\theta}({\tilde{\mathbf{u}}}(k),k,\hat{\mathbf{c}})$ receives as input the noised agent actions $\tilde{\mathbf{u}}(k)\in\mathbb{R}^{A\times T\times 2}$ from the clean actions ${\mathbf{u}}(0)=f^{-1}(\mathbf{x}(0))$, which are computed using an inverse dynamic model $f^{-1}$ from the ground truth states $\mathbf{x}(0)\in\mathbb{R}^{A\times T\times 4}$. Each agent’s action at every timestep $u=[\dot{v},\dot{\psi}]^{T}$ consists of acceleration and yaw rate, while the state comprises coordinates, heading, and velocity ($x,y,\psi,v$). The denoiser also takes in the noise level $k$ and the encoded scene conditions $\hat{\mathbf{c}}$ as inputs. It encompasses several Transformer layers with both self-attention and cross-attention. Initially, the noised actions $\tilde{\mathbf{u}}_{k}$ are converted to noised states $\mathbf{\tilde{x}}(k)=f(\tilde{\mathbf{u}}(k))$, encoded via an MLP and concatenated with the noise level embedding. Then, a self-attention Transformer module is employed to model the joint distribution of future plans across agents. To maintain closed-loop rollout causality, a causal relationship mask (Huang et al., 2023a) is used in the self-attention module, which ensures that information from future timesteps cannot be utilized at the current timestep. Furthermore, a cross-attention Transformer module is used to model the scene-conditional distribution, by relating the noisy trajectories to the encoded scene conditions. This decoding block can be repeated several times, and the final embedding is fed into an MLP decoder to predict the denoised actions ${\hat{\mathbf{u}}}(0)$ and clean states $\mathbf{\hat{x}}(0)$ through a dynamic function $f$.
@@
Behavior Predictor. The behavior predictor $\mathcal{P}_{\psi}$ generates the marginal distributions of possible behaviors for individual agents by directly decoding from the encoded scene conditions $\hat{\mathbf{c}}$. The predictor is composed of $L_{\mathcal{P}}$ cross-attention Transformer decoder layers. To accurately predict the probabilities of possible goals, we utilize type-wise static anchors (end goals) $\{\zeta^{i}\}_{i=1}^{M}$ extracted from data (Shi et al., 2023) as the modality query inputs to the initial Transformer decoder layer. The predictor iteratively refines its predictions through several decoding layers and finally generates $M$ trajectories over the future time for each agent along with their associated scores, represented as $\{\text{Cat}_{M}(\hat{\mathbf{u}}^{a},\hat{\omega}^{a})\}_{a=1}^{A}\in\mathbb{R}^{A\times M\times T\times 5}$. To ensure the kinematic feasibility of these goals, the model predicts action trajectories, which are converted into state trajectories using the same dynamic function $f$. Each waypoint in the state trajectory contains the state ($x,y,\psi,v$) and the probability of the trajectory $\hat{\omega}$.
4.2 4.2 Model Training
@@
which is defined as the the Smooth L1 loss between ground-truth trajectories ${\mathbf{x}}$ and the trajectories $\hat{\mathbf{x}}$ rollout from $\hat{\mathbf{u}}$. Essentially, the denoiser is trained to recover the clean trajectories under various noise levels. At each training step, noise level $k$ is sampled and applied to corrupt the ground-truth trajectories, and the denoiser is optimized to predict the denoised trajectories from the corrupted trajectories. Since the model predicts scene-level joint trajectories, all agent trajectories are affected by the same noise level. The design of the cost function plays a crucial role in generation performance and is later ablated in Section 8.1.
In addition, we utilize a novel log noise schedule defined as:
@@
where $K$ is the maximum diffusion steps and $\delta$ is a scale factor that controls the variance changing rate of the log noise schedule. This schedule ensures that the signal-to-noise ratio (SNR) remains sufficiently low, preventing degradation in inference performance due to short-cut learning. Empirically, we found the performance of this noise schedule is significantly better than the traditional Cosine scheduler (Nichol and Dhariwal, 2021). A more detailed discussion of the noise schedule choice can be found in Section 8.2. The full training procedure of the denoiser is described in Algorithm 1.
Algorithm 1 Training process of denoiser
1:Denoiser $\mathcal{D}_{\theta}$, dataset $D$, denoising steps $K$, dynamics function $f$, inverse dynamics function $f^{-1}$
2:for each training iteration do
3:     ${\mathbf{x}},\mathbf{c}\sim D$triangleright Sample from dataset
4:     Get action trajectory: $\mathbf{u}=f^{-1}({\mathbf{x}})$
5:     $k\sim\mathcal{U}(0,K)$, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$triangleright Sample noise level and Gaussian noise
6:     Add noise to ground-truth: $\tilde{\mathbf{u}}_{k}=\sqrt{\bar{\alpha}_{k}}\mathbf{u}+\sqrt{1-\bar{\alpha}_{k}}\epsilon$
7:     Predict denoised trajectory: $\hat{\mathbf{u}}=\mathcal{D}_{\theta}(\tilde{\mathbf{u}}_{k},k,\mathbf{c},f),\ \hat{{\mathbf{x}}}=f(\hat{\mathbf{u}})$
8:     Compute loss: $\mathcal{L}_{\mathcal{D}_{\theta}}=\mathcal{SL}_{1}(\hat{{\mathbf{x}}}-{\mathbf{x}})$triangleright Use smooth L1 loss
9:     Update denoiser parameters $\theta$
10:end for
@@
$m^{*}=\begin{cases}\arg\min_{i}||\mathbf{ac}^{i}-\mathbf{x}_{T}||,&\text{if }\mathbf{x}_{T}\text{ is valid},\\ \arg\min_{i}||\sum_{t}({\mathbf{\hat{x}}}_{t}^{i}-\mathbf{x}_{t})||,&\text{otherwise},\end{cases}$ (17)
@@
where $\gamma$ is a hyperparameter to balance the importance of tasks.
4.3 4.3 Experimental Setup
System Dynamics. The diffusion model operates in the action space, and we assume that there is a dynamic function that can translate actions to physical states ${\mathbf{x}}_{t+1}=f({\mathbf{x}}_{t},\mathbf{u}_{t})$. We utilize unicycle dynamics as the system dynamics ${\mathbf{x}}_{t+1}=f({\mathbf{x}}_{t},\mathbf{u}_{t})$ to roll out actions into states. This model is shared across all types of agents, including vehicles, pedestrians, and cyclists. The current state of an agent is defined by its global coordinates $(x,y)$, yaw angle $\psi$, and velocities $v_{x},v_{y}$. Given the action of an agent, including acceleration $\dot{v}$ and yaw rate $\dot{\psi}$, and the time length for one step $\Delta t$, the next-step state of the agent is calculated using the following forward dynamics $f$, expressed as:

### v3

@@ III-B Denoising Diffusion Probabilistic Model (DDPM)
where $\mu_{\theta}$ calculates the posterior mean of the noise at k1k-1 step from $\tilde{\mathbf{u}}_{k}$ and DDPM denoiser output $\mathcal{D}_{\theta}(\cdot)$, and $\sigma_{k}$ is the standard deviation according to the fixed noise schedule.
@@
Figure 2: Overview of the proposed VBD model structure. The input scenario context is encoded through a query-centric Transformer scene context encoder. The behavior predictor generates marginal multi-modal trajectories. The denoiser predicts the joint multi-agent future trajectories while attending to themselves and the condition tokens. During inference, the predicted behavior priors or user-defined model-based objectives $\mathcal{J}$ can be used to guide the denoising process to generate desired scenarios.
III-C Controllable and Compositional Generation
We also want to generate scenarios that satisfy a specific user requirement $y$ without retraining the model. For example, $y$ can be defined as the goal or the reference path for individual agents, or it can describe a soft constraint, such as obeying the speed limit or avoiding collisions. We modify the optimization objective to: $\mathcal{J}_{\theta}({\mathbf{x}},\mathbf{u};\mathbf{c})+\mathcal{J}_{y}({\mathbf{x}},\mathbf{u};\mathbf{c})$. Plugging into the EBM representation, we obtain a new conditional distribution: $p(\mathbf{u}|\mathbf{c},y)\propto p(\mathbf{u}|\mathbf{c})p(y|\mathbf{u},\mathbf{c})$, where $p(\mathbf{u}|\mathbf{c})$ is the data distribution we approximated through generative modeling and $p(y|\mathbf{u},\mathbf{c})$ is the likelihood of $y$. This immediately resembles the compositionality in EBM [12] and Classifier Guidance in diffusion models [11]. Specifically, we can sample the reverse process with a conditional score function:
@@ III-C Controllable and Compositional Generation
where $p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})$ is the likelihood of $y$ given the noised action $\tilde{\mathbf{u}}$ at step $k$. It is important to note that $p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})$ is not equivalent to the likelihood of $y$ in the data distribution $p(y|\mathbf{u},\mathbf{c})$, therefore it is typically required to train a separate model [11, 26]. However, we can utilize a practical approximation to the gradient of the noised likelihood with the gradient of $\mathcal{J}_{y}$ for guidance, which enables flexible composition and controllability with additional objectives without training.
In the reverse diffusion process, the guidance is implemented by subtly altering the predicted mean of the model at each denoising step [65, 64]. Initially, we can directly approximate $\log p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})\approx\log p(y|\tilde{\mathbf{u}},\mathbf{c})=\mathcal{J}_{y}({\mathbf{x}}(\tilde{\mathbf{u}}),\tilde{\mathbf{u}};\mathbf{c})$; therefore, the denoising mean is modified to impose guidance in the form of a score function $\mathcal{J}_{y}$ as:
@@
where parameter $\lambda$ controls the strength of the guidance.
However, calculating the gradient based on the mean of noisy actions is difficult and inaccurate, and thus manipulating the noise mean using the noisy gradient can result in errors and instability. To address this, we utilize an alternative approach [27] to approximate $\log p_{k}(y|\tilde{\mathbf{u}},\mathbf{c})\approx\mathcal{J}_{y}({\mathbf{x}}(\mathcal{D}_{\theta}(\tilde{\mathbf{u}})),\mathcal{D}_{\theta}(\tilde{\mathbf{u}});\mathbf{c})$. This means calculating the objective function using the one-step generation result from the denoiser rather than the noisy mean. The modification in the denoising step is expressed as:
@@
where gradient $\nabla_{\tilde{\mathbf{u}}_{k}}$ is calculated with respect to the noisy actions $\tilde{\mathbf{u}}_{k}$ and necessitates differentiation through the denoiser.
IV Method
IV-A Model Structure
@@
The Versatile Behavior Diffusion (VBD) model consists of three main components, as illustrated in Fig. 2. The scene encoder $\mathcal{E}_{\phi}:\mathbf{c}\mapsto\hat{\mathbf{c}}$ encodes the scene context $\mathbf{c}$ into its latent representation $\hat{\mathbf{c}}$ using query-centric attention Transformers [46]. Leveraging rich scene context information from the encoder, the denoiser $\mathcal{D}_{\theta}:(\hat{\mathbf{c}},\tilde{\mathbf{u}},k)\mapsto\hat{\mathbf{u}}$ directly predicts a joint clean control sequence $\hat{\mathbf{u}}$ from $\hat{\mathbf{c}}$ and noised control $\tilde{\mathbf{u}}$ at step $k$. The behavior predictor $\mathcal{P}_{\psi}:(\hat{\mathbf{c}},\{\zeta^{i}\}_{i=1}^{M})\mapsto\{\text{Cat}_{M}(\hat{\mathbf{u}}^{a},\hat{\omega}^{a})\}_{a=1}^{A}$ predicts an $M$-mode marginal categorical trajectory distribution of each agent from $\hat{\mathbf{c}}$ with the help of a set of representative static end-point anchors $\{\zeta^{i}\}_{i=1}^{M}$ extracted from data [46]. All three modules utilize a stack of query-centric self-attention and cross-attention blocks.
@@
Scene Context. The scene conditions are divided into three categories: agents $\mathbf{c}_{a}\in\mathbb{R}^{A\times T_{h}\times D_{a}}$, map polylines $\mathbf{c}_{m,pl}\in\mathbb{R}^{M_{l}\times M_{p}\times D_{p}}$, and traffic lights $\mathbf{c}_{m,tl}\in\mathbb{R}^{M_{t}\times D_{t}}$. Here, $T_{h}$ denotes the number of historical steps, $M_{l}$ the number of polylines, $M_{p}$ the number of waypoints per polyline, $M_{t}$ the number of traffic lights, and $M_{l}+M_{t}=M$ represents the combined count of map elements. The feature sizes for agents, polylines, and traffic lights are represented by $D_{a}$, $D_{p}$, and $D_{t}$, respectively. The agent history tensor records each agent’s historical state, including $x,y$ coordinates, heading angle ($\psi$), velocities ($v_{x},v_{y}$), and bounding box dimensions ($l,w,h$), along with the agent type. Each map polyline, comprising $M_{p}$ waypoints, includes attributes such as $x,y$ coordinates, direction angle, the traffic light state controlling the lane, and lane type. The traffic lights tensor encompasses $x,y$ coordinates of stop points and the state of each traffic light. Before encoding, positional attributes of all elements are converted into their local coordinate systems; for agents, the reference point is their last recorded state, and for map polylines, it is the location of the first waypoint.
@@
Scene Encoder. We encode the agent history tensor utilizing a shared GRU network, which is then combined with the agent type embedding. For map polylines, an MLP is employed for encoding, followed by max-pooling along the waypoint axis; for traffic lights, we encode only their light status using an MLP. These tensors are then concatenated to form the initial scene encoding. The initial scene encoding is further processed using $L_{\mathcal{E}}$ query-centric Transformer layers to symmetrically encode the interrelationships among scene components, resulting in a uniform dimension for scene features $\hat{\mathbf{c}}\in\mathbb{R}^{(A+M)\times D}$. In this approach, each scene element is translated into its local coordinate system and encoded with query-centric features, and the relative position of each pair of scene elements is calculated and encoded as edge attributes.
[…]
@@
Initially, the noised actions $\tilde{\mathbf{u}}_{k}$ are converted to noised states $\mathbf{\tilde{x}}(k)=f(\tilde{\mathbf{u}}(k))$, encoded via an MLP, and concatenated with the noise level embedding. Then, a self-attention Transformer module is employed to model the joint distribution of future plans across agents. To maintain closed-loop rollout causality, a causal relationship mask [22] is used in the self-attention module, ensuring that information from future timesteps cannot be utilized at the current timestep. Furthermore, a cross-attention Transformer module is used to model the scene-conditional distribution by relating the noisy trajectories to the encoded scene conditions. This decoding block can be repeated several times, and the final embedding is fed into an MLP decoder to predict the denoised actions ${\hat{\mathbf{u}}}(0)$ and clean states $\mathbf{\hat{x}}(0)$ through a dynamics function $f$.
@@
Behavior Predictor. The behavior predictor $\mathcal{P}_{\psi}$ generates the marginal distributions of possible behaviors for individual agents by directly decoding from the encoded scene conditions $\hat{\mathbf{c}}$. The predictor is composed of $L_{\mathcal{P}}$ cross-attention Transformer decoder layers. To accurately predict the probabilities of possible goals, we utilize type-wise static anchors (end goals) $\{\zeta^{i}\}_{i=1}^{M}$ extracted from data [46] as the modality query inputs to the initial Transformer decoder layer. The predictor iteratively refines its predictions through several decoding layers and finally generates $M$ trajectories over the future time horizon for each agent along with their associated scores, represented as $\{\text{Cat}_{M}(\hat{\mathbf{u}}^{a},\hat{\omega}^{a})\}_{a=1}^{A}\in\mathbb{R}^{A\times M\times T\times 5}$. To ensure the kinematic feasibility of these goals, the model predicts action trajectories, which are converted into state trajectories using the same dynamics function $f$. Each waypoint in the state trajectory contains the state ($x,y,\psi,v$) and the probability of the trajectory $\hat{\omega}$.
IV-B Model Training
@@
which is defined as the Smooth L1 loss between ground-truth trajectories $\mathbf{x}$} and the trajectories $\hat{\mathbf{x}}$ rolled out from $\hat{\mathbf{u}}$. Essentially, the denoiser is trained to recover the clean trajectories under various noise levels. At each training step, noise level $k$ is sampled and applied to corrupt the ground-truth trajectories, and the denoiser is optimized to predict the denoised trajectories from the corrupted trajectories. Since the model predicts scene-level joint trajectories, all agent trajectories are affected by the same noise level. The design of the cost function plays a crucial role in generation performance.
In addition, we utilize a log noise schedule defined as:
@@
where $K$ is the maximum diffusion step and $\delta$ is a scale factor that controls the variance changing rate of the log noise schedule. This schedule ensures that the signal-to-noise ratio (SNR) remains sufficiently low, preventing degradation in inference performance due to shortcut learning. Empirically, we found the performance of this noise schedule is significantly better than the Cosine noise scheduler [38].
@@
[…]
@@
where $\gamma$ is a hyperparameter to balance different tasks.
V Experimental Setup
V-A System Dynamics
The diffusion model operates in the action space, and we assume that there is a dynamic function that can translate actions to physical states ${\mathbf{x}}_{t+1}=f({\mathbf{x}}_{t},\mathbf{u}_{t})$. We utilize unicycle dynamics as the system dynamics ${\mathbf{x}}_{t+1}=f({\mathbf{x}}_{t},\mathbf{u}_{t})$ to roll out actions into states. This model is shared across all types of agents, including vehicles, pedestrians, and cyclists. The current state of an agent is defined by its global coordinates $(x,y)$, yaw angle $\psi$, and velocities $v_{x},v_{y}$. Given the action of an agent, including acceleration $\dot{v}$ and yaw rate $\dot{\psi}$, and the time length for one step $\Delta t$, the next-step state of the agent is calculated using the following forward dynamics $f$: